Using the expectations hypothesis theory for the term structure of interest rates, determine the expected return for securities with maturities of two, three, and four years based on the following data. (Input your answers as a percent rounded to 2 decimal places.)

Interest Rate
1-year T-bill at beginning of year 1 4 %
1-year T-bill at beginning of year 2 7 %
1-year T-bill at beginning of year 3 6 %
1-year T-bill at beginning of year 4 9 %
Answer #1

Given interest rates for T-Bill,

Interest Rate
1-year T-bill at beginning of year 1 F(0,0,1) 4.00%
1-year T-bill at beginning of year 2 F(0,1,2) 7.00%
1-year T-bill at beginning of year 3 F(0,2,3) 6.00%
1-year T-bill at beginning of year 4 F(0,3,4) 9.00%

So, spot rate for 2 years f(0,2) = ((1+F(0,0,1))*(1+F(0,1,2)))^(1/2) - 1 = (1.04*1.07)^(1/2) - 1 = 5.5%

Spot rate for 3 years f(0,3) = ((1+F(0,0,1))*(1+F(0,1,2))*(1+F(0,2,3)))^(1/3) - 1 = (1.04*1.07*1.06)^(1/3) - 1 = 5.66%

Spot rate for 4 years f(0,4) = ((1+F(0,0,1))*(1+F(0,1,2))*(1+F(0,2,3))*(1+F(0,3,4)))^(1/4) - 1 = (1.04*1.07*1.06*1.09)^(1/4) - 1 = 6.48%